# The low multipoles in the Pantheon+SH0ES data

### v1

@@ 4.3 Bulk + monopole analysis
We also model the data by adding a monopole to the bulk velocity, the parameter $\gamma$ of Eq. (2.18). Interestingly, while adding a quadrupole with its five free parameters reduces the $\chi^{2}$ by the modest value $\Delta\chi_{Q}^{2}\simeq 5.9$ with respect to the analysis including only a bulk velocity, by adding a monopole characterized by just one free parameter, $\gamma$, we gain a $\Delta\chi^{2}_{M}\simeq 6.85$. The inclusion of the monopole also leads to a slight increase of $H_{0}$, by $0.9\sigma$ and to a decrease of $\Omega_{{\rm m}}$ by about $1.2\sigma$, see Table 6. The increase of $H_{0}$ can be understood as follows: a negative value of $\gamma$ leads to an increase in $z_{q}$ hence the measured $d_{L}(z)=\bar{d}_{L}(z_{q})>\bar{d}_{L}(z)$. And since $H_{0}$ is inversely proportional to $\bar{d}_{L}(z)$, this implies a larger $H_{0}$. At $z<0.01$ this reduction of $\bar{d}_{L}(z)$ is about 1.2%, but due to the reduced value of $\Omega_{{\rm m}}$, it decays rapidly and is only about 0.02% at $z=0.5$. The difference $\bar{d}_{L}(z,\Omega_{{\rm m}}=0.316,$ $H_{0}=74.2)-\bar{d}_{L}(z,\Omega_{{\rm m}}=0.34,$ $H_{0}=73.2)$ crosses zero at $z\simeq 1$, above which the Pantheon+ dataset contains no SNIa.
However, like the quadrupole the monopole does not affect the dipole, ${\mathbf{v}}^{(\rm bulk)}$. This again is a consequence of the fact that the different multipoles are orthogonal functions. As the background luminosity distance is a monopole, only the monopole of the peculiar velocity, i.e. its radial component, can affect the it and thereby modify the inferred cosmological parameters.
@@
$\gamma$ [km/s] $|{\mathbf{v}}^{(\rm bulk)}|$[km/s] ra(bulk)[°] dec(bulk)[°] $H_{0}$[km/s/Mpc] $\Omega_{{\rm m}}$
$-106\pm 40$ $318\pm 40$ $197\pm 11$ $-52.8^{+5.0}_{-6.0}$ $74.2\pm 1.1$ $0.316\pm 0.019$
Table 6: Constraints on parameters for the monopole and dipole inferred in the Pantheon+ data set without imposing any redshift cut. For the sake of simplicity in presenting the results, we omit the constraints on dM.
The contour plots for this analysis are shown in Fig. 7. Note the correlation between $\gamma$ and $\Omega_{{\rm m}}$ and the anti-correlation of $\gamma$ and $H_{0}$. Despite the relatively low best fit amplitude of the monopole, about $1/3$ of the bulk velocity, its impact on the cosmological parameters is quite strong. Note also that the monopole is distinguished from the background luminosity distance only via its redshift dependence. Nevertheless, it is detected with a significance of more than $2\sigma$. As we shall see, this is mainly due to its strong effect at very low redshift.
@@ 4.4 The full bulk + quadrupole + monopole analysis
Finally we model the redshift by adding all, the radial velocity (monopole), the bulk velocity (dipole) and the quadrupole. With respect to the analysis allowing only for a bulk velocity (dipole), we gain a $\Delta\chi^{2}\simeq 9.31$ in this model which has six additional parameters. As we have seen in the previous section, most of this improvement is due to the monopole. The contour plots of this analysis are shown in Fig. 8 and the best fit values with $1\sigma$ error bars are reported in Table 7. As already in the pure monopole analysis, the monopole amplitude, $\gamma$ is correlated with $\Omega_{{\rm m}}$ and $H_{0}$. The best fit values of $H_{0}$ and $\Omega_{{\rm m}}$ are affected by the presence of $\gamma$, but the values found in the original Pantheon+ analysis [23] remain consistent within $1\sigma$ with our results.
@@
Figure 8: Contour plots for the full bulk velocity, quadrupole and monopole analysis described in section 4.4 with three different cuts in the redshift of the Supernovae. More cuts are shown in Figs. 15 in Appendix A
@@ 4.5 Applying redshift cuts
We repeat our bulk + monopole + quadrupole analysis to sub-portion of the Pantheon+ dataset obtained removing all the supernovae with a redshift smaller than a certain values $z_{\rm cut}$. We obtain the contour plots reported in Figs. 89 and in Appendix A, Figs. 13 and 14. In Table 8 we report the number of supernovae included within a given redshift cut.
@@
Table 8: Number of supernova lightcurves in each sub-portion of the Pantheon+ dataset obtained removing all the supernovae with a redshift smaller than $z_{\rm cut}$. (Note that while the Pantheon+ compilation contains 1550 different SNIa, it has 1701 lightcurves as several supernovae have been observed in more than one experiment.) In particular, in the last column we provide also the number of SNe in Cepheid-host galaxies in each sub-dataset.
Figure 9: Contour plots for the Pantheon+ data set with three different cuts in the redshift of the Supernovae. More cuts are shown in Figs. 13 and 14 in Appendix A.
@@
Our results are also summarized in tables 7, where the full analysis is presented and in table 9 where we do not include the monopole. Note that already at $z=0.005$, the monopole is no longer detected at more than $1\sigma$. It is significant only at very low redshifts where it is multiplied by a factor $1/H_{0}r(z)\simeq 1/z$ Nevertheless, its presence does somewhat raise the inferred value of the Hubble parameter, albeit within $1\sigma$ and not in the direction which would reduce the Hubble tension. Note also that for the first redshift cut only 9 Supernovae in galaxies without Cepheids and 27 supernovae in galaxies with Cepheids are removed, see table 8. Hence most of the monopole signal comes from modelling these 9 lowest redshift supernovae.
In table 7 we do not include the redshift cut $z_{\rm cut}=0.1$ since for this cut the likelihood contours become very flat and good convergence cannot be reached. For this reason we also include the redshift-cut study without the monopole reported in table 9. This MCMC converges well also for $z_{\rm cut}=0.1$. The results for the bulk velocity and the quadrupole of this analysis are in good agreement with the full analysis. It is interesting to note that while for $z_{\rm cut}\geq 0.0375$ the dipole, i.e., the bulk velocity, is no longer detected and more than $1.5\sigma$, see table 9 and Figs. 13 and 14, the eigenvalues $\lambda_{1}$ and $\lambda_{3}$ of the quadrupole remain non-zero even at 95% confidence. Contrary to the bulk velocities, the eigenvalues $\lambda_{1}$ and $\lambda_{3}$ of the quadrupole are even increasing with redshift cut. This means that above $z_{\rm cut}=0.0375$, corresponding to a distance $R=112h^{-1}$pc, the angular fluctuations in the luminosity distance are better fitted with a quadrupole than with a dipole. This is not so surprising as the quadrupole represents fluctuating velocity field roughly on the scale of the redshift cut while the bulk velocity is assumed to be constant on all scales relevant in the analysis.
@@ 4.6 $\chi^{2}$ analysis
It is also interesting to study the improvement of the fit when including the monopole and quadrupole. As shown in Table 10, this strongly depends on the redshift cut. Introducing no cut the fit is significantly improved and as we have seen this is mainly due to the monopole. At the very low cuts of $z_{\rm cut}=0.005$ and $z_{\rm cut}=0.015$ corresponding to a radius of $15h^{-1}$pc and $30h^{-1}$pc, the improvement is not significant considering that we have introduced 5 or, with monopole 6, additional parameters. However, for $z_{\rm cut}\geq 0.0175$, $\Delta\chi^{2}$ is monotonically increasing and for $z_{\rm cut}\geq 0.025$, the improvement obtained by including a quadrupole (and a at this redshift irrelevant monopole) is, even if not overwhelming, more substantial. This is due to the fact that even for the highest redshift cuts, $\lambda_{1}$ and $\lambda_{3}$ are non-zero within 95% confidence. In the contrary, their best fit values are even increasing. Even though also the error bars increase, the significance of $\lambda_{1}$ and $\lambda_{3}$ simply measured as (best fit value)/error also increases with redshift. This is not the case for ${\mathbf{v}}^{\rm(bulk)}$ which can vanish within less than 90% confidence for $z_{\rm cut}\geq 0.0375$ and becomes even less significant with increasing redshift, see Table 9 and Figs. 13 and 14 in Appendix A.
@@
      $0.1$       $9.52$       -
Table 10: $\Delta\chi^{2}$ differences for different redshift cuts between the best fit dipole determined in our previous analysis [22] and the hypothesis of a bulk motion with only quadrupole correction (second column) and the hypothesis of a bulk motion with both quadrupole and monopole correction (third column).
Finally, we note that, as shown for the dipole in Fig. 2, there is an excellent agreement between the approach where we implement the quadrupole and monopole perturbation as a redshift correction or at the level of the definition of luminosity distance (2.8).
@@ 4.7 Mock tests
Figure 10: Visual comparison of the redshift distribution given by Pantheon+ (orange histogram) and the mock distribution (blue histogram) used in section 4.7
@@
Figure 11: We compare the best fit parameters from the original Pantheon+ data (orange) with the ones from Mock data in which both, the bulk velocity and the quadrupole have been removed.
@@
In order to confirm the significance of the monopole, quadrupole and the bulk velocity, we have also compared the analysis of the true Pantheon+ data with the analysis of a mock dataset created using an artificial redshift distribution (the comparison with the original redshift distribution is in Fig. 10), the distance moduli computed according to Eqs. (2.162.182.19), fixing a constant $|{\mathbf{v}}^{(\rm bulk)}|=400$km/s in direction $({\rm ra,dec})=(100$°, $20$°), so that the monopole and quadrupole vanish, and a cosmology given by the fiducial values $\Omega_{{\rm m}}=0.338$, $H_{0}=73.4$. Moreover, for simplicity, we neglected supernovae hosted in cepheids and fixed dM=0, and we used a diagonal covariance whose elements are the same as the diagonal of the covariance matrix of the Pantheon+ analysis.
@@
For $H_{0}$, $\Omega_{{\rm m}}$ and the bulk velocity we are able to recover the values chosen for the construction of the mock dataset. At the same time, we do not find any monopole or quadrupole as they are detected in the real data. In Fig. 11 we show the results from a parameter estimation of the Mock dataset compared to the real data for dipole and quadrupole only. The bulk velocity inserted in the Mock data and the cosmological parameters used are indicated by dashed vertical and horizontal lines. Clearly, the input parameters are very well reproduced and all three eigenvalues of the quadrupole peak at zero. The same is true for the monopole. This confirms our interpretation that the bulk velocity, the monopole and the quadrupole are present in the data.

### v2

@@ 4.4 Bulk + monopole analysis
We also model the data by adding a monopole to the bulk velocity, the parameter $\gamma$ of eq. (2.18). The unperturbed cosmological $\bar{d}_{L}(z)$ of course also represents a monopole. But the redshift dependence of the velocity monopole, given by $(1+z)^{2}H^{-1}(z)A(z)$ is quite different from $\bar{d}_{L}(z)$ so that this degeneracy is lifted. Note especially that contrary to $\bar{d}_{L}(z)$, the redshift contribution from the velocity monopole does not vanish for $z\rightarrow 0$. This is due to the fact that a radial velocity of the source modifies the measured redshift so that $d_{L}=0$ now no longer coincides with vanishing redshift. This makes the monopole at low redshift very distinct from $\bar{d}_{L}(z)$.
Interestingly, while adding a quadrupole with its five free parameters reduces the $\chi^{2}$ by the modest value $\Delta\chi_{Q}^{2}\simeq 5.9$ with respect to the analysis including only a bulk velocity of section 4.2, by adding a monopole characterized by just one free parameter, $\gamma$, we gain a $\Delta\chi^{2}_{M}\simeq 6.85$. The inclusion of the monopole also leads to a slight increase of $H_{0}$, by $0.9\sigma$ and to a decrease of $\Omega_{{\rm m}}$ by about $1.2\sigma$, see third row of table 4. The increase of $H_{0}$ can be understood as follows: Considering eqs. (2.16) to(2.18) we see that a negative value of $\gamma$ leads to an increase in $z_{q}$ with respect to its value for $\gamma=0$ hence the measured $d_{L}(z)=\bar{d}_{L}(z_{q})>\bar{d}_{L}(z)$. And since $H_{0}$ is inversely proportional to $\bar{d}_{L}(z)$, this implies a larger $H_{0}$. At $z<0.01$ this reduction of $\bar{d}_{L}(z)$ is about 1.2%, but due to the reduced value of $\Omega_{{\rm m}}$, it decays rapidly and is only about 0.02% at $z=0.5$. The difference $\bar{d}_{L}(z,\Omega_{{\rm m}}=0.316,$ $H_{0}=74.2)-\bar{d}_{L}(z,\Omega_{{\rm m}}=0.34,$ $H_{0}=73.2)$ crosses zero at $z\simeq 1$, above which the Pantheon+ dataset contains no SNIa.
However, like the quadrupole the monopole does not affect the dipole, ${\mathbf{v}}^{(\rm bulk)}$. This again is a consequence of the fact that the different multipoles are orthogonal functions. As the background luminosity distance is a monopole, only the monopole of the peculiar velocity, i.e. its radial component, can affect it and thereby modify the inferred cosmological parameters.
@@
Model $\gamma$ $|{\mathbf{v}}_{\odot}|$ ra dec $|{\mathbf{v}}^{(\rm bulk)}|$ ra(bulk) dec(bulk) $H_{0}$ $\Omega_{{\rm m}}$
[km/s] [km/s] [°] [°] [km/s] [°] [°] [km/s/Mpc]
Sec. 4.1 - $318\!\pm\!40$ $140\!\pm\!7.7$ $42^{+7}_{-6}$ $73.2\pm 1.0$ $0.340\pm 0.018$
Sec. 4.2 - - - - $320\pm 40$ $203\pm 11$ $-52.5^{+4.8}_{-5.5}$ $73.2\pm 1.0$ $0.340\pm 0.018$
Sec. 4.4 $-106\!\pm\!40$ $318\pm 40$ $197\pm 11$ $-52.8^{+5.0}_{-6.0}$ $74.2\pm 1.1$ $0.316\pm 0.019$
Table 4: Constraints on parameters for the models we discussed in sections 4.14.2 and 4.4 without imposing any redshift cut. For the sake of simplicity in presenting the results, we omit the constraints on dM. Here and in all the following results tables, the errors show the 68% confidence intervals obtained when analysing the MCMC chains with getdist. They are purely statistical errors, as a consequence, they should be interpreted with care.
The contour plots for this analysis are shown in figure 7. Note the correlation between $\gamma$ and $\Omega_{{\rm m}}$ and the anti-correlation of $\gamma$ and $H_{0}$. Despite the relatively low mean value amplitude of the monopole, about $1/3$ of the bulk velocity, its impact on the cosmological parameters is quite strong. Despite the fact that the monopole is distinguished from the background luminosity distance only via its redshift dependence, it is detected with a significance of more than $2\sigma$. As we shall see, this is mainly due to its strong effect at very low redshift.
@@ 4.5 The full bulk + quadrupole + monopole analysis
Finally we model the redshift by adding all, the radial velocity (monopole), the bulk velocity (dipole) and the quadrupole. With respect to the analysis allowing only for a bulk velocity we performed in section 4.2, we gain a $\Delta\chi^{2}\simeq 9.31$ in this model which has six additional parameters. As we have seen in the previous section, most of this improvement is due to the monopole. The $\Delta\chi^{2}$ for the quadrupole alone is only somewhat larger than 5 which is the expected value for 5 new parameters and not a truly better fit, while for the single parameter $\gamma$, $\chi^{2}$ is reduced by more than 6.
The contour plots of this analysis are shown in figure 9 (green contours and lines) and the mean values with $1\sigma$ error bars are reported in table 5 (first line). As already in the pure monopole analysis, the monopole amplitude, $\gamma$ is correlated with $\Omega_{{\rm m}}$ and $H_{0}$. The mean values of $H_{0}$ and $\Omega_{{\rm m}}$ are affected by the presence of $\gamma$, but the values found in the original Pantheon+ analysis [23] remain consistent within $1\sigma$ with our results.
Figure 8: Visualisation of the constraints on monopole, dipole and quadrupole as reported in table 5.
@@
Figure 9: Contour plots for the full bulk velocity, quadrupole and monopole analysis described in section 4.5 with three different cuts in the redshift of the supernovae. More cuts are shown in figure 19 in Appendix A.
@@ 4.6 Applying redshift cuts
[…]
@@
Table 6: Number of supernova lightcurves in each sub-portion of the Pantheon+ dataset obtained removing all the supernovae with a redshift smaller than $z_{\rm cut}$. (Note that while the Pantheon+ compilation contains 1550 different SNIa, it has 1701 lightcurves as several supernovae have been observed in more than one experiment.) For information, in the last column we provide also the number of SNe in Cepheid-host galaxies in each sub-dataset.
Figure 10: Contour plots for the bulk velocity and monopole analysis described in section 4.4 with three different cuts in the redshift of the supernovae. More cuts are shown in figures 17 and 18 in Appendix A.
@@
Our results are also summarized in table 5, where the full analysis is presented, and in table 7 where we do not include the monopole. Note that already at $z=0.005$, the monopole is no longer detected at more than $1\sigma$. It is significant only at very low redshifts where it is multiplied by a factor $1/H_{0}r(z)\simeq 1/z$. Nevertheless, its presence does somewhat raise the inferred value of the Hubble parameter, albeit within $1\sigma$ and not in the direction which would reduce the Hubble tension. Note also that for the first redshift cut only 9 supernovae in galaxies without Cepheids and 27 supernovae in galaxies with Cepheids are removed, see table 6. Hence most of the monopole signal comes from modelling these 9 lowest redshift supernovae.
This MCMC converges well also for $z_{\rm cut}=0.1$. The results for the bulk velocity and the quadrupole of this analysis are in good agreement with the full analysis. It is interesting to note that while for $z_{\rm cut}\geq 0.0375$ the dipole, i.e., the bulk velocity, is no longer detected at more than $1.5\sigma$, see table 5 and figures 19 and 20, the eigenvalues $\lambda_{1}$ and $\lambda_{3}$ of the quadrupole remain non-zero even at 95% confidence. Contrary to the bulk velocities, the eigenvalues $\lambda_{1}$ and $\lambda_{3}$ of the quadrupole are even increasing with redshift cut. This means that above $z_{\rm cut}=0.0375$, corresponding to a distance $R=112h^{-1}$pc, the angular fluctuations in the luminosity distance are better fitted with a quadrupole than with a dipole. This is not so surprising as the quadrupole represents fluctuating velocity field roughly on the scale of the redshift cut while the bulk velocity is assumed to be constant on all scales relevant in the analysis.
@@ 4.7 $\chi^{2}$ analysis
It is also interesting to study the improvement of the fit when including the monopole and quadrupole. As shown in table 8, this strongly depends on the redshift cut. Introducing no cut the fit is significantly improved and as we have seen this is mainly due to the monopole. At the very low cuts of $z_{\rm cut}=0.005$ and $z_{\rm cut}=0.015$ corresponding to a radius of $15h^{-1}$pc and $30h^{-1}$pc, the improvement is not significant considering that we have introduced 5 or, with monopole 6, additional parameters. However, for $z_{\rm cut}\geq 0.0175$, $\Delta\chi^{2}$ is monotonically increasing and for $z_{\rm cut}\geq 0.025$, the improvement obtained by including a quadrupole (and monopole which is irrelevant at this redshift) is, even if not overwhelming, more substantial. This is due to the fact that even for the highest redshift cuts, $\lambda_{1}$ and $\lambda_{3}$ are non-zero within 95% confidence. On the contrary, their mean values are even increasing. Despite the fact that also the error bars increase, the significance of $\lambda_{1}$ and $\lambda_{3}$ simply measured as (mean value)/error also increases with redshift. This is not the case for ${\mathbf{v}}^{\rm(bulk)}$ which can vanish within less than 90% confidence for $z_{\rm cut}\geq 0.0375$ and becomes even less significant with increasing redshift, see table 7 and figures 17 and 18 in Appendix A. We finally note that whenever tested we found a slightly smaller $\chi^{2}$ when the correction is applied as a redshift correction, eq. (2.19), than when it is applied on the distance $d_{L}$ directly, eq. (2.8).
@@
      $0.1$       $9.52$       $9.80$
Table 8: $\Delta\chi^{2}$ differences for different redshift cuts between the mean value dipole determined in our previous analysis [22] and the hypothesis of a bulk motion with only quadrupole correction (second column) and the hypothesis of a bulk motion with both quadrupole and monopole correction (third column).
@@ 4.8 Mock tests
Figure 11: Visual comparison of the redshift distribution given by Pantheon+ (orange histogram) and the mock distribution (blue histogram) used in section 4.8.
@@
Figure 12: We compare the mean value parameters from the original Pantheon+ data (orange) with the ones from Mock data in which both, the monopole and the quadrupole have been removed. Note that $\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$ together with $\lambda_{1}+\lambda_{2}+\lambda_{3}=0$ enforces $\lambda_{1}\leq 0$ and $\lambda_{3}\geq 0$.
@@
In order to confirm the significance of the monopole, quadrupole and the bulk velocity (dipole in the supernova data), we have also compared the analysis of the true Pantheon+ data with the analysis of a mock dataset created using an artificial redshift distribution (the comparison with the original redshift distribution is shown in figure 11), the distance moduli computed according to eqs. (2.162.182.19), fixing a constant $|{\mathbf{v}}^{(\rm bulk)}|=400$km/s in direction $({\rm ra,dec})=(100$°, $20$°), so that the monopole and quadrupole vanish, and a cosmology given by the fiducial values $\Omega_{{\rm m}}=0.334$, $H_{0}=73.6$. Moreover, for simplicity, we neglected supernovae hosted in cepheids, fixing dM=0, and we used a diagonal covariance whose elements are the same as the diagonal of the covariance matrix of the Pantheon+ analysis.
@@
For $H_{0}$, $\Omega_{{\rm m}}$ and the bulk velocity we are able to recover the values chosen for the construction of the mock dataset. At the same time, we do not find any monopole or quadrupole, while they are detected in the real data. In figure 12 we show the results from a parameter estimation of the Mock dataset compared to the real data for dipole and quadrupole only. The bulk velocity inserted in the Mock data and the cosmological parameters used are indicated by dashed vertical and horizontal lines. Clearly, the input parameters are very well reproduced and all three eigenvalues of the quadrupole are well consistent with zero. The same is true for the monopole. This confirms our interpretation that the bulk velocity, the monopole and the quadrupole are really present in the data.
5 Systematics checks
We tested wether our signal was dominated by a single supernova. For doing so, we first computed the $\chi^{2}$ differences for the full bulk + monopole + quadrupole analysis between considering all the supernovae and removing one supernova at a time. We plot the differences as function of the redshift in fig. 13. We then run an MCMC removing the 4 supernovae with redshift $z\leq 0.1$ contributing the most to $\chi^{2}$ (i.e. the supernovae whose $\Delta\chi^{2}\geq 10$). We considered only $z<0.1$ supernovae since they are the ones mainly constraining the multipoles. From the MCMC run, we obtained essentially identical contours as in fig. 9, proving that the our signal is not dominated by a one or a few supernovae.
Figure 13: $\chi^{2}$ differences as function of redshift for the full bulk + monopole + quadrupole analysis between considering all the supernovae and removing one supernova at a time.
[…]
$\Delta v^{\text{(bulk)}}=7{\rm km/s}\,,\qquad({\rm\Delta ra,\Delta dec})=(32,-1.7)\,,$ (5.1)
while the difference between the monopole mean values is
$\Delta\gamma=-89{\rm km/s}$ (5.2)
While the monopole becomes even more negative and actually more significant, the trend for the bulk velocity is less clear. We therefore just add the difference due to Milky Way dust extinction as systematic uncertainties in quadrature to the errors inferred from the Pantheon+ covariance matrix that is given in table 5. The resulting errors for the bulk velocity are shown in table 9.
$v^{(\text{bulk})}$[km/sec] ra $[^{o}]$ dec $[^{o}]$
$323\pm 50.5$ $199\pm 34$ $-51.2^{+5.4}_{-7.2}$
Table 9: The bulk velocity when adding a systematic error due to Milky Way dust extinction.
Clearly, adding these possible errors does not change our main results. The bulk velocity remains significant and even though the right ascension is now within $1\sigma$ of the observer velocity $v_{\odot}$ given in (1.1), the declination is still very far away, it rather becomes even more negative.
Figure 14: Comparison of the full supernovae distribution (orange histogram) with the distribution of the 1320 supernovae with MWEBV0.05absent0.05$\leq$ 0.05≤ 0.05 (blue histogram).
Figure 15: Contour plots for the full bulk velocity, quadrupole and monopole analysis described in section 4.5 considering all the supernovae (orange contours) and the 1320 supernovae with MWEBV0.05absent0.05$\leq$ 0.05≤ 0.05 (blue contours).